Mechanics

The potential energy of the simple harmonic oscillator when the particle is half way up to its end point is? If the total energy is E?
a. E/4 b. E/2 c. 2E/3 d.E/8

Regina Phalange
7 Years agoGrade

1 Answer

Arun
7 Years ago

P.E. = (1/2) mω²y² ⇒ At y = a/2 ⇒ (1/2) (mω²a²/4)
Total energy (E) = P.E. at extreme position = (1/2) mω²a²
P.E. = (1/4).[(1/2) mω²a²] = E/4